OpenStudy (anonymous):

The sum of the digits of a two digit number is 12. The value of the numberis equal to 11 times the tens digit. What is the product of its digits?

OpenStudy (anonymous):

let the number be xy sum of digits = 12 therefore x+y = 12 value of the number= 10*(number in the tens place) + 1*(number in the units place) = 10x+y, which is given to be 11 times the tens digits. Therefore, 10x + y = 11x so you need to solve the two equations and find out x and y and finally just multiply x*y

OpenStudy (anonymous):

is the answer "6" ?

OpenStudy (anonymous):

x=6 and y=6 so the number is 66. u just need to find the product now

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

yes. Keep it up. try solving more such questions

OpenStudy (anonymous):

thank you. actually it's my daughter's problems. she's too tired to do her work, that is why i'm here to ask for help.

OpenStudy (anonymous):

:D thats fine.